OpenStudy (anonymous):

The electrical resistance r of a wire varies directly with its length l and inversely with the square of its diameter d. If 200 m of wire with diameter 6 mm has a resistance of 16 ohms, what is the resistance of 300 m of wire with a diameter of 8 mm?

OpenStudy (anonymous):

r is the resistance k is the constant l is the length d is the diameter. r = k*l r = (1/k)*(d^2)

OpenStudy (anonymous):

Thank you so much! I really appreciate it. I was lost. Lol.

OpenStudy (anonymous):

your welcome.